b+c, c+a, a+b are in H.P., then a^2, b^2, c^2will be in


a. AP


b. GP


c. HP


d. None of these


 


Explain the steps and how to approach in such questions

Grade:12

4 Answers

Aman Bansal
12 years ago

Dear Saawan,

The correct answer is b. GP

BEST OF LUCK..!!!!

Plz approve the answer....!!

We are all IITians and here to help you in your IIT JEE preparation.

Now you can win exciting gifts by answering the questions on Discussion Forum. So help discuss any query on askiitians forum and become an Elite Expert League askiitian.

Now you score 5+15 POINTS by uploading your Pic and Downloading the Askiitians Toolbar : Click here to download the toolbar.. 

AMAN BANSAL

SAAWAN Pawan VERMA
12 years ago

i know that the answer is b.

But tell me the explaination.!!

Pavani Bandi
12 years ago

b+c, c+a, a+b are in H.P

1/b+c, 1/c+a, 1/a+b are in A.P.

2/c+a=1/b+c+1/a+b

by solving u get

2b2 =a2+c2

So a2,b2,care in A.P.

Pavani Bandi
12 years ago

I think answer is A.P.